# boxparampadding=1.5pt, padding-left=2pt, padding-right=2pt, height=7.5pt, background-color=lightgray, border-radius=2pt, border-right-width=1pt, border-bottom-width=1pt, border-color=gray ExploreSelf: Fostering User-driven Exploration and Reflection on Personal Challenges with Adaptive Guidance by Large Language Models

### v2

@@ 5.2. Safety Considerations
Although our target participant group did not specifically include individuals experiencing mental illness, we prepared safety protocols for the study based on the work of O’Leary et al. (O’Leary et al., 2018), whose study involved peer support chat amongst individuals experiencing mental illness. This was a mindful procedure considering that ExploreSelf encourages participants to disclose personal issues that might be sensitive and potentially traumatic, and participants could have mental health issues that were not disclosed to the experimenter.
@@ 5.3. Study Setup and Procedure
Briefing. The experimenter shared her screen with presentation slides (refer to our supplementary material) and described the motivation and goal of our study. Here, participants were fully informed that the goal of this study is not about their specific personal challenges but to evaluate the design and user experience of ExploreSelf to explore and reflect on those challenges.
@@
Exploration. After the tutorial, the experimenter shared a link to access ExploreSelf and invited the participant to a Zoom breakout room, a virtual space where the participant could stay alone and focus. The experimenter was able to monitor the real-time interaction logs on ExploreSelf to observe the activity patterns. The experimenter and the participant were connected via an instant messenger to communicate about any issues the participant may have during the exploration including help requests. The participant then freely used ExploreSelf to explore their challenge. There was no minimum time requirement, and participants could finish the session whenever they felt they had sufficiently used it, for up to 45 minutes. At the 45-minute mark, the researcher checked in with participants if they needed more time. If so, they were allowed up to an additional 15 minutes, with a maximum session time of 1 hour. Since participants did not have free practice in the tutorial phase, the system offered on-demand help popups on major interface components such as the Theme Selector dialog. To assess the impact of the exploration on participants’ attitudes towards their challenges, we asked participants to fill out the Adult State Hope Scale (Snyder et al., 1996) before and after the exploration phase.
@@
Debriefing. At the end of the session, we removed the breakout room and conducted a semi-structured interview, which took about 30 minutes. Before starting the interviews, we confirmed their willingness to proceed. We asked participants about their experiences with ExploreSelf, their rationales for adding and moving between themes and questions, receptiveness to AI-driven guidance, and how the interface helped or hindered their exploration and reflection.
5.4. Data Analysis
We employed both quantitative and qualitative methods to analyze our data, in order to gain a comprehensive understanding of participant engagement and interaction with ExploreSelf. First, to understand what types of personal challenges participants would explore on our system, participants’ initial narratives were categorized through an iterative open-ended coding process conducted by two authors. The sub-categories were multi-coded to the initial narratives and grouped into broader categories.
To analyze their overall usage and interaction patterns, We collected interaction logs to track key behaviors, such as the number of themes and questions selected by participants, the number of interactions with keywords and comments, and phase-switching behaviors (e.g. between exploration and overview phase). These logs were analyzed to compute descriptive statistics, including the minimum, maximum, mean, and standard deviation for user engagement across different features of the system. Additionally, participants rated their perceived agency both immediately before and after the study by completing the Pathways Subscale of the State Hope Scale (Snyder et al., 1996), using the validated Korean version (Choi et al., 2008). This 4-item measure asks participants to rate statements such as “There are lots of ways around any problem that I am facing now” on an 8-point scale, ranging from 1 (Definitely false) to 8 (Definitely true). The Pathways Subscale of the State Hope Scale, used in this study, captures individuals’ perceptions of their ability to achieve their goals (Snyder et al., 1996; Schleider et al., 2020; Buchlmayer et al., 2024). Scores range from 4 to 32, with higher scores reflecting a stronger belief in the availability of pathways to their goals Appendix B.
We further analyzed debriefing transcripts by conducting thematic analysis (Braun and Clarke, 2006) to identify key themes regarding participants’ experiences with ExploreSelf. Through collaborative discussions, we reviewed, refined, and consolidated the recurring themes until a consensus was achieved on key themes. By combining quantitative metrics from interaction logs and qualitative insights from debriefing interviews, we were able to explore how participants utilized ExploreSelf’s features and how the system supported their reflective practices.
6. Results
6.1. Overall System Usage
@@
Personal interactions and societal
influences affecting connections with others. |
Family conflicts[P1, P8, P14, P15], parenting[P1, P7], social relationships[P6, P8], romantic relationships[P3, P16], societal stigma[P4], and societal pressure[P1, P11, P13, P19] |
@@
Academic stress[P5], job dissatisfaction[P10], time-management concern, goal fulfillment, career uncertainties[P1, P5, P6, P9, P12, P14, P15, P17] |
Self-identity &
character (8 participants) |
@@
Aging[P8, P14], sexual identity[P3], existential concern[P8, P16], self-esteem issue[P2, P7, P8, P13, P19] |
@@
Financial crisis[P8, P9, P11], financial instability[P5, P9, P18], life transitions[P16] |
@@
Managing anxiety[P16, P18], trauma[P4, P8], emotional struggles[P2], physical health concerns[P8] |
@@
Syllables of
initial narrative |
233.63 | 299.98 | 233 | 478 | 39 | 215 | 313 | 445 | 42 | 1347 | 100 | 180 | 198 | 13 | 137 | 1,534 | 25 | 34 | 111 | 239 | 127 |
Total
syllables of responses |
745.47 | 349.65 | 1,261 | 568 | 699 | 388 | 1,273 | 1,019 | 541 | 1,278 | 571 | 325 | 381 | 496 | 1,381 | 417 | 824 | 508 | 864 | 465 | 905 |
@@
Total # of
questions |
11.47 | 7.28 | 17 | 7 | 18 | 3 | 11 | 4 | 15 | 6 | 5 | 4 | 8 | 28 | 12 | 9 | 10 | 12 | 22 | 4 | 23 |
Total # of
revealed keywords |
11.79 | 8.69 | 0 | 9 | 0 | 27 | 6 | 11 | 29 | 18 | 8 | 9 | 24 | 15 | 4 | 18 | 20 | 4 | 6 | 9 | 7 |
Total # of
comment requests |
6.47 | 8.26 | 0 | 1 | 0 | 2 | 0 | 5 | 18 | 4 | 1 | 0 | 6 | 1 | 11 | 3 | 10 | 27 | 9 | 1 | 24 |
@@ 6.2. Sense of Agency over Exploration
In debriefing, participants expressed varying perceptions of their agency while using ExploreSelf, with several factors influencing their sense of control. Some participants felt that the acknowledgment of their wordings in the AI-generated text played a crucial role in shaping the experience. P11 observed that while the AI sometimes paraphrased his inputs into different expressions in its generation, this gave him the impression that the AI was taking an active role. However, he still felt a sense of agency because the AI-generated texts were based on his wordings, highlighting that his expressions ultimately drove the interaction. P8 echoed this sentiment, noting that the system’s ability to suggest themes based on her input made her feel that she was leading the process.
@@
In general, participants appreciated the way ExploreSelf visualized and dynamically structured the breakdowns of their thoughts into a coherent landscape. 10 participants highlighted how this experience differed from their self-journaling practices. P18, for instance, shared that when reflecting on her own, her thoughts would often “spiral in a negative direction” with no clear way to shift perspectives. In contrast, the system allowed her to “step back” and revisit previous thoughts easily, helping her avoid getting stuck in a mental loop. Similarly, P7 remarked that if she had tried to write about her stress and various life challenges on her own, her thoughts would have “jumped from one issue to another,” leading to a scattered reflection. However, with the AI’s structure, she could “stick to one theme” at a time, making the process more focused and less overwhelming.
@@ 6.3. User-directed Navigation
Participants navigated through themes and questions by making deliberate choices about whether to explore deeper into a particular theme (i.e., add a new question to the thread) or move on to a new theme. Some participants managed their emotional engagement with a theme by deciding when to stop drilling into certain topics to avoid emotional distress or over-involvement. P17 explained “while certain questions seemed helpful, continuing to engage with them might have deepened my emotional engagement, but I wasn’t quite ready at the moment. I thought I would come back at some point.” Based on the findings from participant debriefing interviews, we present how participants interacted with the themes and questions in the pathways of processing their thoughts and emotions.
@@ 6.3.1. Expanding Viewpoints through Themes
Participants also emphasized that the suggestions of alternative wordings for the same theme allowed them to engage in their narrative from multiple angles, broadening their perspectives and deepening their reflections. This functionality encouraged participants to think about the same theme in different ways, as seen when P14 explored the theme of “retirement” from multiple viewpoints, such as “anxiety about retirement, purpose, and direction of life after retirement, and feeling unprepared for old age”. These varied expressions also provided flexibility in how participants responded to the themes, prompting different sets of questions depending on how a theme was framed, and also expanded their capacity for self-expression. As P9 noted, the system’s varied prompts allowed him to “respond more creatively and diversely” by offering new ways to articulate their thoughts.
Participants usually made the decision to move on to a new theme when they felt they had gained enough insight or a good understanding of a particular issue. For example, P13 shifted her focus toward solutions once she understood that her low self-esteem was the root of her challenge: “I decided to think about how to address it, rather than keep going deeper into why my self-esteem was low.”
@@ 6.3.2. Organizing Angles and Attitudes through Questions
Participants also valued questions that challenged their usual thinking patterns, prompting reflections on aspects they had never considered before. P2 noted, “I usually set personal goals and feel frustrated when I don’t meet them. The AI asked, ‘Why do you set such standards?’ which made me select it because it questioned something I had always taken for granted, like the fact that the Earth is round. It was an idea I had never thought about, and that’s why it stood out to me.”
In contrast, some participants avoided more abstract or challenging questions, opting for those that “felt easier to answer based on concrete experiences (P5).” To balance the cognitive load, several participants voluntarily alternated between engaging with challenging, thought-provoking questions and those they found more manageable. For example, P5 noted, “It wasn’t easy to think about, but it [ExploreSelf] asked ‘what my strengths are in a competitive setting,’ so I went ahead with [that question] because I figured it would be useful to learn about,” expressing willingness to confront difficult questions if he believed the outcome would ultimately be beneficial to him. Similarly, P8 shared that after writing about difficult and negative experiences, she felt emotionally drained. When presented with multiple follow-up questions, she opted for one that could “shift their thinking in a new direction”, selecting “How would your life change if your financial situation were stable?” over other equally “relevant and effective questions” such as, “What impact is your financial difficulty having on your life?” or “What steps have you taken to resolve your financial challenges?”. This choice allowed her to regulate her emotional state while still maintaining engagement without becoming overwhelmed by the difficult emotions.
@@ 6.3.3. Objectification and Reiteration through AI Summary
Participants generally viewed the AI summary to bring clarity and objectivity to their often complex and unorganized thoughts by distilling the fragmented responses into an organized narrative. In turn, the AI summary seemed to foster iterative reflection; twelve out of 19 participants (63%) revisited the Exploration page back, explored more, and went back to the AI summary page to check the new summary reflecting additional responses (See dark blue bars in Figure 4). P12 remarked, “I liked the summary, but I was curious how it might turn out if I revisited it, so I kept checking. After seeing the summary again, I thought of things I had missed and wrote them down, then went back to see how the updated summary looked. This back-and-forth really helped me organize my thoughts.” This iterative process, where participants reviewed and refined their reflections, enhanced their engagement and allowed for further self-exploration.
@@ 6.4.2. Iteratively Developing and Refining Answers through Comments
In sum, the comments helped participants in three ways. First, the comments provided them with serendipitous insights through a more detailed message. For example, P16 shared that the AI suggested they reflect on physical reactions when reflecting on emotional experience they had written, a perspective she had never considered. Second, participants thought the comment was particularly insightful when it acknowledged with their feelings or thoughts with clear rationales. For instance, P7 noted, “When I was asked about the most meaningful moment I could think of, I responded that while I felt hopeful, I also felt despair because it was a past that would never return. After I clicked on new comment, the AI said something like, ‘It’s natural to feel this way because meaningful moments are often retrospective.’ The fact that it acknowledged my feelings as natural was really meaningful to me.” Lastly, encouraging comments often motivated participants for a continued engagement, though not all experiences were positive. P15 recounted an instance where, after responding to the question “Have you tried any small actions or changes to rekindle a recent passion?” with “There’s nothing I remember.” ExploreSelf gave the comment, “It’s fine to start by thinking this way; take your time to reflect,” and it inspired her to keep thinking and eventually led to a more carefully-thought-out answer. However, not all feedback was well-received: P16 felt that some comments were overly upbeat and didn’t match the gravity of her situation, causing a sense of emotional disconnection.
6.5. Perceived Utility of ExploreSelf
To gauge the utility of ExploreSelf beyond the study, we asked participants about their willingness to try ExploreSelf in the future. All but one participant (18 out of 19; 95%) said that they would like to use ExploreSelf after the study to process the challenges that would arise in the future. One participant chose not to use the tool because he had expected an AI to offer actionable solutions for his challenges. Others envisioned revisiting ExploreSelf when they have recurring struggles “like a tangled ball of yarn” (P9).
@@ 7. Discussion
In this section, we discuss lessons learned from the design and evaluation of ExploreSelf as well as the limitations and future work.
7.1. Fostering User-Driven Exploration and Reflection with Adaptive Guidance by LLMs
In our study, we explored how providing people with structured yet flexible interactions with adaptive guidance provided by LLMs could facilitate exploration and reflection on personal challenges. Starting from the initial narrative, the system offered various forms of exploration components and adaptive guidance, including themes, questions, keywords, comments, and summaries. These components allowed participants to navigate their narratives interactively, managing the process at multiple touchpoints, such as selecting themes, selecting questions, switching between themes, and moving between the exploration and summary phases.
Notably, participants reflected not only while they were articulating answers in response to the questions but also continuously while they were interacting with the system’s adaptive elements, particularly through the selection of both themes and questions. When choosing themes, participants engaged in reflection by examining which themes emerged from their input, assessing which felt most significant, and deciding whether they were ready to explore certain themes or preferred to revisit them later. Similarly, selecting questions within these themes prompted further reflection in a broader cognitive and emotional landscape. Participants evaluated the different pathways their responses could take based on the questions presented, considering how each question aligned with their current emotional and cognitive state.
The interaction with keywords and comments that supported participants to respond to the Socratic questions further enriched their reflections. By anchoring their reflections in specific expressions, keywords helped participants articulate complex emotions and explore perspectives, allowing them to further interpretations, or reject suggested keywords when they felt they did not align with their personal context or perspective. Also, the interaction with the comments, whether it was an encouragement, insight, acknowledgment, or sub-question for iteration, the expectation of receiving feedback after each response encouraged users to remain involved in the reflective process.
In addition, switching between the summary and exploration phases encouraged reflection by providing a more holistic perspective on their progress. By reviewing the LLM-generated summary, participants could step back and assess how their thoughts had evolved, prompting them to revisit the exploration or decide whether to consolidate their insights. The combination of selecting themes, questions, and reviewing summaries created an ongoing cycle of reflection that extended beyond the direct, one-off act of answering a prompt, enriching the overall reflective experience.
The natural language generation capabilities of LLMs played a key role in making the interaction components reflective. Since the wording of the themes, questions, and other guidance were generated by LLMs given participant responses, the elements often directly referred to the participants’ wording rather than using general, abstract text. We suspect that such detailed references to participant’s language made participants feel empathized and engaged, as similarly has been observed in other studies involving conversational agents (e.g., (Wei et al., 2024; Kim et al., 2024a; Jo et al., 2024; Jain et al., 2018)).
As such, our study suggests that designing user interfaces that can visually support flexible and dynamic navigation of thoughts and feelings, as well as provide adaptive guidance powered by generative models, can altogether shape an environment where users can meaningfully process their personal challenges at their own pace, on their own terms.
7.2. Enhancing User Agency of Interaction and Reflection
Our survey results indicate the exploration with ExploreSelf has increased the perceived agency of participants. Also, the participants reported several factors that impacted their perceived agency during reflection, revealing two types of agency at work. First, ExploreSelf enhanced the user agency of interaction, the perception that participants had control over the exploration while interacting with an AI. Notably, participants felt more control when the AI-generated wordings were aligned with or referring to their previous inputs. On the contrary, participants felt the AI was taking the initiative when those wordings were not aligned with their narrative. Sometimes, participants were willing to give control to an AI when they believed that it provided serendipitous insights even with unfamiliar or uncomfortable themes and questions.
On the other hand, we found ExploreSelf also improved user perceived agency in reflection, the perception that participants felt having control or being in charge of their situation. As users felt more self-aware as they engaged with ExploreSelf, they felt more clarity in making choices on which themes to choose, whether they wanted more depth, and which questions to answer. Moreover, the agency that they felt from interaction also helped this process. Navigating themes and keywords, in turn provided them with an increased self-understanding, and the insights brought them to a higher level of situational awareness and their resources to resolve or approach the problem at hand.
We believe this observation that LLM-driven adaptive guidance can promote different layers of user agency in exploring and reflecting on personal challenges carries an important implication in many contexts of human-AI collaboration. Previous work has mostly invested in designing intelligent interfaces that can support user agency in interaction, noting that users would like to take control (Heer, 2019; Moruzzi and Margarido, 2024; Hwang and Won, 2022) in the interaction process and outcome. Our work suggests another type of user agency, at a cognitive level, that is in play while users are interacting with the AI. We find that users want to be in the state of feeling control at face of their personal challenges, for the very lack of it was why they were weighing on their minds. Future work should investigate how AI could be designed in intelligent systems that can unobtrusively encourage the path to enhanced perceived agency in problem exploration.

### v3

@@ 5.2. Safety Considerations
Although our target participant group did not specifically include individuals experiencing mental illness, we prepared safety protocols for the study based on the work of O’Leary et al. (O’Leary et al., 2018), which involved peer support chat amongst individuals experiencing mental illness. This was a mindful procedure considering that ExploreSelf encourages participants to disclose personal issues that might be sensitive and potentially traumatic, and participants could have mental health issues that were not disclosed to the experimenter.
@@ 5.3. Study Setup and Procedure
Briefing. The experimenter shared her screen with presentation slides (refer to our supplementary material) and described the motivation and goal of our study. To encourage participants to authentically engage with the system, we informed them that the goal of this study is not to evaluate or address their specific personal challenges but to understand how our tool helps their exploration and reflection on those challenges.
@@
Exploration. After the tutorial, the experimenter shared a link to access ExploreSelf and invited the participant to a Zoom breakout room, a virtual space where the participant could stay alone and focus. The experimenter was able to monitor the real-time interaction logs on ExploreSelf to observe the activity patterns. The experimenter and the participant were connected via an instant messenger to communicate about any issues the participant may have during the exploration including help requests. The participant then freely used ExploreSelf to explore their challenge. There was no minimum time requirement, and participants could finish the session whenever they felt they had sufficiently used it, for up to 45 minutes. At the 45-minute mark, the researcher checked in with participants if they needed more time. If so, they were allowed up to an additional 15 minutes, with a maximum session time of 1 hour. Since participants did not have free practice in the tutorial phase, the system offered on-demand help popups on major interface components such as the Theme Selector dialog. To assess the impact of the exploration on participants’ attitudes towards their challenges, we asked participants to rate their perceived agency by completing the validated Korean version (Choi et al., 2008) of the Pathways Subscale of the State Hope Scale (Snyder et al., 1996), before and after of this phase. This 4-item measure asks participants to rate statements such as “There are lots of ways around any problem that I am facing now” on an 8-point scale, ranging from 1 (Definitely false) to 8 (Definitely true). The Pathways Subscale of the State Hope Scale (refer to Appendix C), used in this study, captures individuals’ perceptions of their ability to achieve their goals (Snyder et al., 1996; Schleider et al., 2020; Buchlmayer et al., 2024). Scores range from 4 to 32, with higher scores reflecting a stronger belief in the availability of pathways to their goals.
@@
Debriefing. At the end of the session, we removed the breakout room and conducted a semi-structured interview for about 30 minutes. Before starting the interviews, we confirmed their willingness to proceed. We asked participants about their experiences with ExploreSelf, their rationales for adding and moving between themes and questions, receptiveness to AI-driven guidance, and how the interface helped or hindered their exploration and reflection.
@@
Personal interactions and societal influences
affecting connections with others. |
Family conflicts[P1, P8, P14, P15], parenting[P1, P7], social relationships[P6, P8], romantic relationships[P3, P16], societal stigma[P4],
societal pressure[P1, P11, P13, P19] |
@@
Academic stress[P5], job dissatisfaction[P10], time-management concern,
goal fulfillment, career uncertainties[P1, P5, P6, P9, P12, P14, P15, P17] |
Self-identity & character
(8 participants) |
@@
Aging[P8, P14], sexual identity[P3], existential concern[P8, P16],
self-esteem issue[P2, P7, P8, P13, P19] |
@@
Financial crisis[P8, P9, P11], financial instability[P5, P9, P18],
life transitions[P16] |
@@
Managing anxiety[P16, P18], trauma[P4, P8], emotional struggles[P2],
physical health concerns[P8] |
5.4. Data Analysis
We employed both quantitative and qualitative methods to analyze our data, in order to gain a comprehensive understanding of participant engagement and interaction with ExploreSelf. First, to understand what types of personal challenges participants would explore on our system, participants’ initial narratives were categorized through an iterative open-ended coding process conducted by two authors. The first author, who has two years of experience in HCI and mental health research and administered the user study, first open-coded the initial narratives. As the initial narratives usually described multiple aspects, we multi-coded the narratives. Another researcher with a PhD in computer science and 12 years of experience in HCI and health informatics research participated in multiple rounds of discussion to finalize a codebook for sub-category labels. The full research team then iterated multiple sessions of discussions to group the subcategories into five broader categories. Although the coding was primarily bottom-up and grounded from the data, some category labels were informed by Sharma et al. (Sharma et al., 2023)’s definitions of personal issues (e.g., parenting, trauma) and the life values set used in Acceptance and Commitment Therapy (Hayes and Pierson, 2005) (e.g., social relationships). The categorization is reported in Table 3.
To analyze their overall usage and interaction patterns, We collected interaction logs to track key behaviors, such as the number of themes and questions selected by participants, the number of interactions with keywords and comments, and phase-switching behaviors (e.g. between exploration and overview phase). These logs were analyzed to compute descriptive statistics, including the minimum, maximum, mean, and standard deviation for user engagement across different system features.
The Zoom sessions for debriefing interviews were recorded and transcribed later. We qualitatively analyzed the debriefing transcripts using thematic analysis (Braun and Clarke, 2006) to identify key themes regarding participants’ experiences with ExploreSelf. The first author open-coded the transcripts, going through several rounds of iterations. The full research team then discussed and identified patterns and themes through multiple rounds of peer debriefing. From this coding, we surfaced the key themes around the participants’ sense of agency while using ExploreSelf and how the system features supported the exploration, reflection, and self-expression.
6. Results
6.1. Overall System Usage
@@
| Syllables of initial narrative | 233.63 | 299.98 | 233 | 478 | 39 | 215 | 313 | 445 | 42 | 1347 | 100 | 180 | 198 | 13 | 137 | 1,534 | 25 | 34 | 111 | 239 | 127 |
| Total syllables of responses | 745.47 | 349.65 | 1,261 | 568 | 699 | 388 | 1,273 | 1,019 | 541 | 1,278 | 571 | 325 | 381 | 496 | 1,381 | 417 | 824 | 508 | 864 | 465 | 905 |
@@
| Total # of questions | 11.47 | 7.28 | 17 | 7 | 18 | 3 | 11 | 4 | 15 | 6 | 5 | 4 | 8 | 28 | 12 | 9 | 10 | 12 | 22 | 4 | 23 |
| Total # of revealed keywords | 11.79 | 8.69 | 0 | 9 | 0 | 27 | 6 | 11 | 29 | 18 | 8 | 9 | 24 | 15 | 4 | 18 | 20 | 4 | 6 | 9 | 7 |
| Total # of comment requests | 6.47 | 8.26 | 0 | 1 | 0 | 2 | 0 | 5 | 18 | 4 | 1 | 0 | 6 | 1 | 11 | 3 | 10 | 27 | 9 | 1 | 24 |
@@ 6.2. Sense of Agency over Exploration
In debriefing, participants expressed varying perceptions of their agency while using ExploreSelf, with several factors influencing their sense of control. Some participants felt that the acknowledgment of their wordings in the AI-generated text played a crucial role in shaping the experience. P11 observed that while the AI sometimes paraphrased his inputs into different expressions in its generation, this paraphrasing gave him the impression that the AI was taking an active role. However, he still felt a sense of agency because the AI-generated texts were grounded in his original wording, highlighting that his expressions ultimately drove the interaction. P8 echoed this sentiment, noting that the system’s ability to suggest themes based on her input made her feel that she was leading the process.
@@
In general, participants appreciated the way ExploreSelf visualized and dynamically structured the breakdowns of their thoughts into a coherent landscape. Ten participants highlighted how this experience differed from their self-journaling practices. P18, for instance, shared that when reflecting on her own, her thoughts would often “spiral in a negative direction” with no clear way to shift perspectives. In contrast, the system allowed her to “step back” and revisit previous thoughts easily, helping her avoid getting stuck in a mental loop. Similarly, P7 remarked that if she had tried to write about her stress and various life challenges on her own, her thoughts would have “jumped from one issue to another,” leading to a scattered reflection. However, with the AI’s structure, she could “stick to one theme” at a time, making the process more focused and less overwhelming.
@@ 6.3. User-directed Navigation
Participants navigated through themes and questions by making deliberate choices about whether to explore deeper into a particular theme (i.e., add a new question to the thread) or move on to a new theme. Based on the findings from debriefing interviews, we present how participants interacted with the themes and questions in the pathways of processing their thoughts and emotions.
@@ 6.3.1. Expanding Viewpoints through Themes
Participants also emphasized that the suggestions of alternative wordings for the same theme allowed them to engage in their narrative from multiple angles, broadening their perspectives and deepening their reflections. This functionality encouraged participants to think about the same theme in different ways, as seen when P14 explored the theme of ‘retirement’ from multiple viewpoints, such as “anxiety about retirement,” “purpose and direction of life after retirement,” and “feeling unprepared for old age.” These varied expressions also provided flexibility in how participants responded to the themes, prompting different sets of questions depending on how a theme was framed, and also expanded their capacity for self-expression. As P9 noted, the system’s varied prompts allowed him to “respond more creatively and diversely” by offering new ways to articulate their thoughts.
Participants tended to move on to a new theme when they felt they had gained enough insight or a good understanding of a particular issue. For example, P13 shifted her focus toward solutions once she understood that her low self-esteem was the root of her challenge: “I decided to think about how to address it, rather than keep going deeper into why my self-esteem was low.”
@@ 6.3.2. Organizing Angles and Attitudes through Questions
Participants also valued questions that challenged their usual thinking patterns, prompting reflections on aspects they had never considered before. P2 noted, “I usually set personal goals and feel frustrated when I don’t meet them. The AI asked, ‘Why do you set such standards?’ which made me select it because it questioned something I had always taken for granted, like the fact that the Earth is round. It was an idea I had never thought about, and that’s why it stood out to me.” In contrast, some participants avoided more abstract or challenging questions, opting for those that “felt easier to answer based on concrete experiences (P5).” To balance the cognitive load, several participants voluntarily alternated between engaging with challenging, thought-provoking questions and those they found more manageable.
@@ 6.3.3. Objectification and Reiteration through AI Summary
Participants generally viewed the AI summary to bring clarity and objectivity to their often complex and unorganized thoughts by distilling the fragmented responses into an organized narrative. In turn, the AI summary seemed to foster iterative reflection; twelve out of 19 participants (63%) revisited the Exploration page, explored more, and went back to the AI summary page to check the new summary reflecting additional responses (See dark blue bars in Figure 6). P12 remarked, “I liked the summary, but I was curious how it might turn out if I revisited it, so I kept checking. After seeing the summary again, I thought of things I had missed and wrote them down, then went back to see how the updated summary looked. This back-and-forth really helped me organize my thoughts.” This iterative process, where participants reviewed and refined their reflections, enhanced their engagement and allowed for further self-exploration.
@@ 6.4.2. Iteratively Developing and Refining Answers through Comments
Comments generally provided participants with serendipitous insights through a more detailed message, and participants thought a comment was particularly insightful when it acknowledged their feelings or thoughts with clear rationales: “[…] The fact that it acknowledged my feelings as natural was really meaningful to me (P7).” Although the encouraging tone of comments often engaged participants in the exploration, not all experiences were positive. For example, P16 felt that some comments were overly upbeat and didn’t match the gravity of her situation, causing a sense of emotional disconnection.
@@ 7. Discussion
In this section, we discuss lessons learned from the design and implementation of ExploreSelf as well as from the user study. We also reflect on the implications for better supporting LLM-guided self-reflection in a long-term and inclusive setting.
7.1. Supporting User-driven Exploration for Mental Well-being
ExploreSelf facilitated user control over their reflective journey by offering multiple options and layers of adaptive support that encouraged users to navigate their thoughts and emotions. This scaffolded structure allowed participants to tailor their exploration pathways, balancing their cognitive and emotional burdens while facing various aspects of their personal challenges. For example, participants chose to engage more deeply with a theme through Socratic questions or pivoted to new themes when they felt ready, demonstrating varied patterns of exploration (see Figure 7). Consequently, participants showed a significant increase in perceived agency after engaging with ExploreSelf, underscoring its potential to facilitate for users to manage their reflective journey.
Our main design goal with ExploreSelf was to prioritize user agency without directing content, while providing guidance. This goal resulted in the emphasis on exploration, leading to design choices that are different from prior LLM-assisted reflection tools, such as DiaryMate (Kim et al., 2024b) and MindfulDiary (Kim et al., 2024a). For example, DiaryMate employed the concept of writing assistant (Lee et al., 2024a) and used LLMs to generate diary sentences given user inputs like keywords and the model temperature (Kim et al., 2024b). While this approach helped participants complete well-written journal entries, the quality and content of LLM generations would largely influence their reflection and even sometimes interrupt their own stream of thought (Kim et al., 2024b). MindfulDiary incorporated a conversational agent, loosely mimicking a counseling interview with users (Kim et al., 2024a). Although the conversational approach effectively demonstrated care and empathy, the LLM-driven agent mostly held the initiative of conversation and the turn-taking interaction focused the user’s attention on a single question presented at each turn (Kim et al., 2024a). ExploreSelf differs from previous work in that, unlike DiaryMate, it offers users choices about what to explore (i.e., themes and questions) rather than generating narratives on their behalf. Additionally, users could determine their level of engagement with each topic through an interface designed to minimize single-path conversational affordances, as in MindfulDiary. This approach encouraged users to ‘contemplate’ their exploration, enabling user-directed control over the process.
In this work, we explored a novel design approach to invite user agency throughout a self-paced exploration for mental well-being. ExploreSelf targeted lay individuals who may not require clinical diagnoses but any of those who would willingly seek assistance in navigating personal challenges. In this context, its approach follows the principles of psychotherapy, particularly its emphasis on client autonomy (Scheel, 2011). Guided by self-determination theory (Deci and Ryan, 2013), it is widely agreed that supporting clients in autonomously exploring and sustaining a process of change is crucial for a successful outcome (Ryan and Deci, 2008; Scheel, 2011; Patall et al., 2008). Similarly, ExploreSelf provided an environment where users can process their challenges at their own pace and on their own terms, which mirrors the role of therapist—fostering a space where individuals feel empowered to explore their emotions, identify goals, and develop strategies for personal growth (Ryan et al., 2011)—illustrating the potential of designing adaptive guidance with LLMs.